# Local EGOP for Continuous Index Learning

### v2

@@ B.3.7 Noisy Manifold
Thus the rates of Lemma 28 is achieved. Plugging these rates into our risk bounds yields $\sqrt{\det\Sigma_{i}}=\Theta(t_{i}^{(2d+1)/4})$, $W(\mu_{t})=O(t_{i})$, thus yielding the desired rate.
Appendix C Extended Discussion
The most obvious problem left open in our approach is the question of accurate EGOP estimation. One path to achieve this is to show that function gradients can also be estimated accurately up to the prescribed error rates in the domains of interest. At initialization, this is of course reasonable, as we focus on a small isotropic region for which the analysis is standard, and an estimator can achieve a fixed level of bias at an $O(n^{-1/2})$ rate. However, in the iteration that immediately follows this can no longer be expected. As a canonical setting, consider the second order anisotropy $g^{T}\Sigma_{i}g=\Theta(t_{i}),v^{T}\Sigma_{i}v=\Theta(\sqrt{t_{i}})$, $v\perp g$. What must be verified is how the elongation impacts pointwise gradient estimates.
When regressing on such an oblique region, the estimated gradient corresponds to the average of the gradients on the ellipsoid. Thus, even if the anisotropy induces bias, it is reasonable to expect that the AGOP, the aggregate of the resulting estimated gradient outerproducts, is very similar to the EGOP. Indeed, we anticipate that the EGOP off-gradient components are on the order of $\Theta(\sqrt{t_{i}})$, the same magnitude as the anticipated bias correction. The supervised noisy manifold hypothesis does not fundamentally change this calculation, as the additional newly introduced non-decaying components of $\Sigma_{i}$ correspond exactly to the directions where the gradient remains constant, and thus present a benign factor in gradient estimation.

### v3

@@ B.3.7 Noisy Manifold
Thus the rates of Lemma 28 is achieved. Plugging these rates into our risk bounds yields $\sqrt{\det\Sigma_{t}}=\Theta(h_{t}^{(2d+1)/4})$, $W(\mu_{t})=O(h_{t})$, thus yielding the desired rate. ∎
Appendix C EGOP Approximation
In this section, we show that metric anisotropy not only improves point regression, but accelerates the estimation of the EGOP matrix in our two model settings. We argue that the smoothed gradients, while potentially biased estimators at individual points, in the aggregate form a matrix that differs from the population quantity by a benign perturbation, leading to an equivalent convergence analysis at adequate sample sizes.
Define the smoothed gradient and EGOP as
$\nabla_{\Sigma}f(x)=\underset{v}{\operatorname{argmin}}\min_{a}\int\|f(z)-a-v(z-x)\|^{2}\,dN(x,\Sigma),\quad\mathcal{L}_{\Sigma}(\mu)=\int\nabla_{\Sigma}f\nabla_{\Sigma}f^{T}\,d\mu.$
We study the covariance recurrence
$\Sigma_{t+1}=h_{t+1}[\beta\hat{\mathcal{L}}_{\Sigma_{t}}(\mu_{t})+(1-\beta)\hat{\mathcal{L}}_{\Sigma_{t-1}}(\mu_{t-1})]^{-1}=:h_{t+1}[\hat{M}_{t}]^{-1}.$ (9)
Lemma 29 (Smoothed Gradient).
There exists $T_{x},C$, $T_{x}(\Sigma)\leq C\|\Sigma\|$ such that, for $\mu=N(x^{*},\Sigma),$
$\nabla_{\Sigma}f=\nabla f+T_{x}(\Sigma),\quad\mathcal{L}_{\Sigma}(\mu)=\mathcal{L}(\mu)+\mathbb{E}_{\mu}[gT_{X}(\Sigma)^{T}+T_{X}(\Sigma)g^{T}]+O(\|\Sigma\|^{2}).$
Proof.
The standard OLS solution yields
$\nabla_{\Sigma}f(x)=\Sigma^{-1}\operatorname{Cov}_{N(x,\Sigma)}(X,f(X))=\mathbb{E}_{N(x,\Sigma)}[\nabla f(X)],$
with the last equality following from Stein’s Identity (Stein, 1981). Applying a Taylor expansion, we see that
$\mathbb{E}_{N(x,\Sigma)}[\nabla f(X)]=\int\nabla f(x)+H(y-x)+T_{x}(y-x,y-x)\,dN(0,\Sigma)=:\nabla f(x)+T_{x}(\Sigma),$
where $T_{x}$ is a second order remainder. For the second claim, the proof follows by a simple expansion,
$\mathcal{L}_{\Sigma}(\mu)$ $=\int\nabla_{\Sigma}f\nabla_{\Sigma}f^{T}\,d\mu$
$=\int\nabla f\nabla f^{T}\,d\mu+\int T_{x}(\Sigma)\nabla f^{T}+\nabla fT_{x}(\Sigma)\,d\mu$
$=\mathcal{L}(\mu)+\int T_{x}(\Sigma)g^{T}+gT_{x}(\Sigma)\,d\mu++\int T_{x}(\Sigma)(\nabla f-g)^{T}+(\nabla f-g)T_{x}(\Sigma)\,d\mu,$
satisfying the claim, with full details being similar to those in the proof of Lemma 3. ∎
This expansion is satisfactory for the full-rank setting, indeed, one can substitute $\mathcal{L}_{\Sigma}$ into the recurrence of Equation 2, and arrive at the same expansion of Lemma 3. A more refined expansion is necessary for the continuous index setting. Let $\operatorname{Null}_{x}$ denote the null subspace at point $x$.
Lemma 30.
(Anisotropic Smoothed Gradient) Let $(X,Y)$ satisfy the supervised noisy manifold hypothesis, $D>2d$. There exists $T_{x,\Sigma_{\operatorname{Null}_{x}}}$ and a uniform constant $C$ satisfying $T_{x,\Sigma_{\operatorname{Null}_{x}}}(A)\leq C\|A\|$, such that
$\nabla_{\Sigma}f(x)=\nabla f(x)+T_{x,\Sigma_{\operatorname{Null}_{x}}}(\Sigma_{\operatorname{Null}_{x}^{\perp}}-[\pi_{\operatorname{Null}_{x}^{\perp}}\Sigma\pi_{\operatorname{Null}_{x}}]\Sigma_{\operatorname{Null}_{x}}^{-1}[\pi_{\operatorname{Null}_{x}}\Sigma\pi_{\operatorname{Null}_{x}^{\perp}}])$
Proof.
We combine Stein’s identity with Lemma 1 to yield
$\nabla_{\Sigma}f(x)=\mathbb{E}_{N(x,\Sigma)}[\nabla f(X)]$ $=\int\nabla f(y)\,dN(y;x,\Sigma)$
$=\int\left(\int\nabla f(y)\,d\mu_{v}\right)\,d\nu(v)=\nabla f(x)+\int T_{v}(\Sigma_{v})d\nu(v)$
$=:\nabla f(x)+T_{x,\Sigma_{\operatorname{Null}_{x}}}(\Sigma_{\operatorname{Null}_{x}^{\perp}}-[\pi_{\operatorname{Null}_{x}^{\perp}}\Sigma\pi_{\operatorname{Null}_{x}}]\Sigma_{\operatorname{Null}_{x}}^{-1}[\pi_{\operatorname{Null}_{x}}\Sigma\pi_{\operatorname{Null}_{x}^{\perp}}]).$
Lemma 31 (Anisotropic Rotation).
For $x$ within a neighborhoood of $x^{*}$,
$\|\pi_{\operatorname{Null}_{x}^{\perp}}\Sigma\pi_{\operatorname{Null}_{x}}\|=O(\|\pi_{\operatorname{Null}^{\perp}_{x^{*}}}[x-x^{*}]\|+\|\pi_{\operatorname{Null}^{\perp}}\Sigma\pi_{\operatorname{Null}}\|),\quad\|\Sigma_{\operatorname{Null}^{\perp}_{x}}^{-1}\|\leq 2/\lambda_{\min}(\Sigma_{\operatorname{Null}^{\perp}}),$
$\|\Sigma_{\operatorname{Null}^{\perp}_{x}}\|=O(\|\pi_{\operatorname{Null}^{\perp}_{x^{*}}}[x-x^{*}]\|^{2}+\|\Sigma_{\operatorname{Null}^{\perp}}\|+\|\pi_{\operatorname{Null}^{\perp}_{x^{*}}}[x-x^{*}]\|\|\Sigma\pi_{\operatorname{Null}^{\perp}}\|).$
Proof.
We first show that
$\|\pi_{\operatorname{Null}_{x}}-\pi_{\operatorname{Null}_{x^{*}}}\|=O(\pi_{\operatorname{Null}_{x^{*}}^{\perp}}[x-x^{*}]),$
from which the desired result follows easily. As characterized in the proof of Lemma 5, $\operatorname{Null}_{x}=\phi_{\pi(x),\nabla f(x)}(\mathcal{T}_{x})^{\perp}$, for $\phi_{\pi(x),u}(w):=\mathrm{I\!I}_{\pi(x)}(u,w)$. Thus we see that this subspace changes smoothly in the displacement of the basepoint $\pi(x)$ or the gradient $\nabla f(x)$, i.e. along $\operatorname{Null}^{\perp}_{x}$.
Thus, setting $D_{x}:=\pi_{\operatorname{Null}_{x}^{\perp}}-\pi_{\operatorname{Null}_{x^{*}}^{\perp}}$,
$\|\pi_{\operatorname{Null}_{x}^{\perp}}\Sigma\pi_{\operatorname{Null}_{x}}\|$ $=\|(\pi_{\operatorname{Null}_{x^{*}}^{\perp}}+D_{x})\Sigma(\pi_{\operatorname{Null}_{x^{*}}}-D_{x})\|$
$\leq\|\pi_{\operatorname{Null}^{\perp}}\Sigma\pi_{\operatorname{Null}}\|+C\|D_{x}\|+C^{\prime}\|D_{x}\|^{2}$
$=O(\|\pi_{\operatorname{Null}^{\perp}}\Sigma\pi_{\operatorname{Null}}\|+\|\pi_{\operatorname{Null}^{\perp}_{x^{*}}}[x-x^{*}]\|),$
as desired. The third claim follows in the exact same way, stopping at the second line of the derivation. The second claim follows by continuity. ∎
It is again necessary to localize and extend $f$ so that these constraints are globally active, or contract the covariance cap $\zeta_{t}\to 0$ at a logarithmic rate.
Lemma 32 (Anisotropic Gradient Bound).
Let $(X,Y)$ satisfy the supervised noisy manifold hypothesis, $D>2d$. There exists $T_{x,\Sigma_{\operatorname{Null}_{x}}}$ and a uniform constant $C$ satisfying $T_{x,\Sigma_{\operatorname{Null}_{x}}}(A)\leq C\|A\|$, such that
$\nabla_{\Sigma}f(x)=\nabla f(x)+T_{x,\Sigma_{\operatorname{Null}_{x}}}(\Sigma_{\operatorname{Null}^{\perp}})+R(x),$
for $R(x)=O(\|\pi_{\operatorname{Null}^{\perp}}[x-x^{*}]\|^{2}+\|\pi_{\operatorname{Null}^{\perp}}[x-x^{*}]\|\|\pi_{\operatorname{Null}^{\perp}}\Sigma\|+\|\pi_{\operatorname{Null}^{\perp}}\Sigma\|^{2}+\|\Sigma_{\operatorname{Null}^{\perp}}\|)$.
Proof.
By Lemma 30 we can express
$\nabla_{\Sigma}f(x)$ $=\nabla f(x)+T_{x,\Sigma_{\operatorname{Null}_{x}}}(\Sigma_{\operatorname{Null}_{x}^{\perp}}-[\pi_{\operatorname{Null}_{x}^{\perp}}\Sigma\pi_{\operatorname{Null}_{x}}]\Sigma_{\operatorname{Null}_{x}}^{-1}[\pi_{\operatorname{Null}_{x}}\Sigma\pi_{\operatorname{Null}_{x}^{\perp}}])$
$=:\nabla f(x)+T_{x,\Sigma_{\operatorname{Null}_{x}}}(\Sigma_{\operatorname{Null}^{\perp}})+R(x).$
That $R(x)$ satisfies the appropriate bound is an immediate consequence Lemma 31 and the Lipschitz continuity of $T$. ∎
Lemma 33 (Anisotropic Smoothed EGOP).
Let $(X,Y)$ satisfy the supervised noisy manifold hypothesis, $D>2d$. Then
$\mathcal{L}_{\Sigma}(\mu)=\mathcal{L}(\mu)+\mathbb{E}_{\mu}[gT_{X,\Sigma_{\operatorname{Null}_{X}}}(\Sigma_{\operatorname{Null}^{\perp}})^{T}+T_{X,\Sigma_{\operatorname{Null}_{X}}}(\Sigma_{\operatorname{Null}^{\perp}})g^{T}]+R(\Sigma),$
for
$R(\Sigma)=O(\|\Sigma_{\operatorname{Null}^{\perp}}\|+\|\pi_{\operatorname{Null}^{\perp}}\Sigma\|^{2}+\sqrt{\|\Sigma_{\operatorname{Null}^{\perp}}\|}\|\pi_{\operatorname{Null}^{\perp}}\Sigma\|).$
Proof.
This result is an immediate consequence of Lemma 32 in combination with Isserlis’s Theorem (Isserlis, 1918), see Lemma 3 for a similar computation. ∎
Thus we have achieved a remainder control that is tight under anisotropy, where $\Sigma$ may stretch in the $\operatorname{Null}$ direction without severe consequence. Our primary assumption will be regarding the accuracy of gradient estimation.
Lemma 34 (EGOP Approximation).
Suppose that $\|\nabla_{\Sigma}\hat{f}-\nabla_{\Sigma}f\|_{L^{2}(\mu)}=O(n^{-1}\det(\Sigma_{t}^{-1/2}))$. Then,
$\hat{\mathcal{L}}_{\Sigma}(\mu)=\mathcal{L}_{\Sigma}(\mu)+O\left(\frac{1}{n\sqrt{\det(\Sigma_{t})}}\right).$
Proof.
We compute,
$\|\hat{\mathcal{L}}_{\Sigma}(\mu)-\mathcal{L}_{\Sigma}(\mu)\|$ $=\left\|\int\nabla_{\Sigma}\hat{f}\nabla_{\Sigma}\hat{f}^{T}-\nabla_{\Sigma}f\nabla_{\Sigma}f^{T}\,d\mu\right\|$
$\leq\int\|\nabla_{\Sigma}\hat{f}-\nabla_{\Sigma}f\|\|\nabla_{\Sigma}f\|+\|\nabla_{\Sigma}\hat{f}-\nabla_{\Sigma}f\|\|\nabla_{\Sigma}\hat{f}\|\,d\mu$
$=O\left(\frac{1}{n\sqrt{\det(\Sigma_{t})}}\right).$
We note that the indicated rate is typical when utilizing standard nonparametric estimators, such as via local linear regression, and can be achieved by taking the covariance cap $\zeta\to 0$ at an appropriate rate. This condition is less easily satisfied for fully anisotropic $\Sigma_{t}$.
Proposition 4 (Full Rank Empirical Guarantee).
Let $H$ be full rank, and suppose that $\|\nabla_{\Sigma_{t}}\hat{f}-\nabla_{\Sigma_{t}}f\|_{L^{2}(\mu_{t})}\leq Cn^{-1}\det(\Sigma_{t}^{-1/2})$. Adopt the assumptions of Theorem 1. Then, for $\Sigma_{t}$ satisfying Equation 9, for $t_{n}=\frac{4\log n}{(D+5)\log(1/r)}$,
$\mathbb{E}\left[\int(\hat{P}_{\hat{M}_{t_{n}}}(f)-f)^{2}\,d\mu_{t_{n}}\right]=O\left(n^{-\frac{4}{D+5}}\right).$
Proof.
By Lemma 29 we see that $\hat{\mathcal{L}}(\mu)$ satisfies an equivalent expansion to Lemma 3 up to $O(n^{-1}\det(\Sigma_{t}^{-1/2}))$ approximation error. As argued in the proof of Theorem 1, the second order recurrence is asymptotically autonomous so long as the empirical error is negligible, satisfying $n^{-1}\det(\Sigma_{t}^{-1/2})=O(h_{t})$. Evaluating,
$n^{-1}\det(\Sigma_{t_{n}}^{-1/2})=n^{-1}h_{t_{n}}^{-(d+1)/4}=O(h_{t_{n}}).$
Proposition 5 (Continuous Index Empirical Guarantee).
Let $(X,Y)$ satisfy the supervised noisy manifold hypothesis, and $\operatorname{rank}(H)=2d$. Suppose that $\|\nabla_{\Sigma_{t}}\hat{f}-\nabla_{\Sigma_{t}}f\|_{L^{2}(\mu_{t})}\leq Cn^{-1}\det(\Sigma_{t}^{-1/2})$. Adopt the assumptions of Theorem 3. Then, for the recurrence
$\Sigma_{t}=h_{t+1}[\beta\hat{\mathcal{L}}_{\Sigma}(\mu_{t})+(1-\beta)\hat{\mathcal{L}}_{\Sigma}(\mu_{t-1})]^{-1},$
selecting $t_{n}=\frac{4\log n}{(2d+5)\log(1/r)}$,
$\mathbb{E}\left[\int(\hat{P}_{M_{t_{n}}}(f)-f)^{2}\,d\mu_{t_{n}}\right]=O\left(n^{-\frac{4}{2d+5}}\right).$
Proof.
The proof again follows immediately from our previous analysis, exploiting the robustness to perturbations we already verified. First however, we must verify that the expansions of Lemma 33 and Lemma 26 are equivalent, i.e. we must verify that
$\|\pi_{\operatorname{Null}^{\perp}}\Sigma_{t}\|^{2},\sqrt{\|\Sigma_{t,\operatorname{Null}^{\perp}}\|}\|\pi_{\operatorname{Null}^{\perp}}\Sigma_{t}\|=O(\|\Sigma_{t,\operatorname{Null}^{\perp}}\|),$ (10)
along the smoothed recurrence. We do this first neglecting noise. The base case is the same, and we adopt the inductive hypothesis in the proof of Theorem 3. Condition (10) is weaker than this inductive hypothesis, and the remainder of the argument is identical. Thus, it suffices to check that empirical error is negligible, satisfying $n^{-1}\det(\Sigma_{t}^{-1/2})=O(h_{t})$. Evaluating,
$n^{-1}\det(\Sigma_{t_{n}}^{-1/2})=n^{-1}h_{t_{n}}^{-(2d+1)/4}=O(h_{t_{n}}).$
C.1 Additional Settings of Interest
Example 5 (Spheres).
$H=\begin{bmatrix}T&b\\ b^{T}&0\end{bmatrix}$
Figure 8: Eigenvalue decay of $\Sigma_{t}$ with Local EGOP Learning applied to data satisfying the noisy manifold hypothesis about $S^{2}$. The orthogonal exhibits light decay, approaching second order anisotropy asymptotically.